# How does online shopping affect offline price sensitivity?

### v1

@@ 8.2.3 Alternative Control Groups
While we do our best to account for differences in observable characteristics between adopters and non-adopters in §$\S$8.2.2, it may be that there are certain time-specific unobserved factors that only affect adopters or would-be adopters, or disproportionately affect them more than non-adopters, thus potentially making non-adopters an unsuitable control group. We conduct an additional robustness check, leveraging the fact that we have multiple cohorts of consumers who adopted online shopping at different times. Following Narang and Shankar (2019) and Manchanda et al. (2015), we consider an alternative specification where later adopter cohorts serve as control groups for earlier adopter cohorts. Since all of these cohorts adopt online shopping at some point, the differences in any time-specific unobservables (that correlate with factors that drive the decision of whether a consumer chooses to adopt online shopping or not) should be minimized across these different adoption cohorts, at least relative to non-adopters. The results from this alternative CS estimator are shown in Web Appendix §$\S$§E.5. Overall, we see that the conclusions remain similar to those from our main specification, i.e., that consumers become more offline price-sensitive after adopting online shopping.
9 Implications for Retailer Pricing and Profitability
So far, we have established that when consumers begin to shop online, their price sensitivity in the offline channel increases. We now examine how the retailer should modify its offline pricing strategy in response to these customer-level changes. In particular, given the firm’s current level of online adoption, how much offline profit can be gained by properly accounting for the change in price elasticity relative to a scenario where the firm ignores this effect? To quantify the profit implications, we conduct counterfactual analyses that compare offline profits under pricing policies that incorporate the online adoption effect vs. those that do not. In §$\S$9.1, we first outline how a naive retailer who ignores the online adoption effect would (incorrectly) estimate the offline price elasticities. Next, in §9.2, we characterize the profit-maximizing prices that would be set under the true demand model (accounting for the online adoption effect) and under the naive demand model (ignoring the online adoption effect) for two pricing strategies: uniform pricing and personalized pricing. In §9.3, we describe how we compute the profits under four potential counterfactual pricing strategies and quantify the profit implications of accounting for the online adoption effect. We also mention some caveats of our counterfactual results.
9.1 Naive demand model without adoption effect
First, we characterize how a naive retailer would estimate consumers’ offline price sensitivities. In this case, we assume that the retailer uses a demand model that does not account for online adoption effects.
Following notation of the demand model in Equation (7) in §6, we specify the mean utility for consumer $i$ choosing brand $j$ at purchase occasion $t$ in this case as follows:
$\tilde{\psi}_{ijt}=\tilde{\alpha}_{0}\textrm{Price}_{jt}+\tilde{\alpha}_{1}\mathcal{I}\{\textrm{BuyLastMonth}_{ijt}\}+\tilde{\beta}_{1}\textrm{CF\_Price}_{jt}+\tilde{\delta}_{ij}~{}~{}\forall j\in\{1,\ldots,J\}\$ (20)
Here, $\tilde{\alpha}_{0}$ is the naive estimate of the price sensitivity parameter. $\tilde{\alpha}_{1}$ and $\tilde{\beta}_{1}$ are coefficients for the state dependence variable and price control function, respectively, and $\tilde{\delta}_{ij}$ denotes customer-brand fixed effects. The corresponding choice probability for consumer $i$ choosing brand $j$ on purchase occasion $t$ is then given by $\tilde{Pr}_{ijt}=\frac{\exp(\tilde{\psi}_{ijt})}{1+\sum_{j^{\prime}=1}^{J}\exp(\tilde{\psi}_{ij^{\prime}t})}$. We follow the estimation procedure described in §6.3 and report the parameter estimates in Table A20 in Web Appendix F. Compared to the parameter estimates in Table 6, we find that the estimated price coefficient $\tilde{\alpha}_{0}$ from the naive model reflects a weighted average of price sensitivities for both adopters and non-adopters. This naive estimate serves as a key input for computing the counterfactual profit-maximizing prices that the retailer would set if they ignored the effect of online adoption.
9.2 Optimal Offline Price Characterization Under Counterfactual Scenarios
We now characterize the profit-maximizing offline prices for a given month, under two pricing strategies: (1) Uniform pricing, where all consumers who visit an offline store during a particular time period see the same price for a given brand, and (2) Personalized pricing, where each consumer may potentially receive a different price vector depending on their demographics, prior purchase behavior, and under our demand model, their online shopping status (whether they have adopted online shopping and how much of their cumulative spend is on the online channel). In practice, the personalized pricing strategy would need to be implemented using targeted coupons (Rossi et al. 1996). Personalized pricing is being increasingly adopted by firms in retail settings and has been the focus of recent research in marketing (Dubé and Misra 2023; Smith et al. 2023; Jain et al. 2024). Therefore the impact of the online adoption effect on both uniform and personalized pricing is of interest for managers.
Let $\mathcal{M}_{k}$ be a generic demand model, where the subscript $k$ denotes the model used. Specifically, $\mathcal{M}_{N}$ denotes the Naive model defined in Equation (20), whereas $\mathcal{M}_{O}$ refers to Our model in Equation (7). We now characterize the optimal offline pricing for a generic demand model $\mathcal{M}_{k}$ and then derive the profit-maximizing price vectors for both the uniform and personalized pricing strategies.
Let $N$ be the market size, which is the number of offline shoppers who make at least one offline store visit during the month. Since we focus on monthly profit, we drop the month index ($t$) for notational simplicity. Let $\mathbf{p}\in\mathcal{R}^{J\times 1}$ denote the price vector, with $p_{j}$ being the price of brand $j$. We use the notation $p_{j}$ for price here (instead of Pricej) since price is now a decision variable rather than an observable. Under the naive model $\mathcal{M}_{N}$, the mean utility for consumer $i$ choosing brand $j$ based on Equation (20) can be written as:
$\psi_{ij}(\mathbf{p};\mathcal{M}_{N})=\tilde{\alpha}_{0}p_{j}+\tilde{\alpha}_{1}\mathcal{I}\{\textrm{BuyLastMonth}_{ij}\}+\tilde{\beta}_{1}\textrm{CF\_Price}_{j}+\tilde{\delta}_{ij}$
In contrast, the mean utility of our demand model $\mathcal{M}_{O}$ based on Equation (7) can be written as:
$\psi_{ij}(\mathbf{p};\mathcal{M}_{O})$ $=\alpha_{0}p_{j}+\alpha_{1}\textrm{cumPercOnlineSpend}_{i}+\alpha_{2}p_{j}\times\textrm{cumPercOnlineSpend}_{i}$
$+\alpha_{3}\mathcal{I}\{\textrm{BuyLastMonth}_{ij}\}+\lambda_{1}\textrm{CF\_Price}_{j}+\lambda_{2}\textrm{CF\_Price\_cumPercOnlineSpend}_{ij}+\delta_{ij}$
9.2.1 Uniform Pricing
Under uniform pricing, the retailer maximizes the expected profit per customer, defined as:
$R(\mathbf{p};\mathcal{M}_{k})$ $=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{J}(p_{j}-m_{j})\times Pr_{ij}(\mathbf{p};\mathcal{M}_{k})=\frac{1}{N}\sum_{i=1}^{N}R_{i}(\mathbf{p};\mathcal{M}_{k})$ (21)
where $m_{j}$ is the cost of brand $j$. $Pr_{ij}(\mathbf{p};\mathcal{M}_{k})=\frac{\exp(\psi_{ij}(\mathbf{p};{\mathcal{M}_{k}})}{1+\sum_{j^{\prime}=1}^{J}\exp(\psi_{ij^{\prime}}(\mathbf{p};{\mathcal{M}_{k}}))}$ refers to the choice probability of consumer $i$ for brand $j$ given price vector $\mathbf{p}$ and demand model $\mathcal{M}_{k}$, and $R_{i}(\mathbf{p};\mathcal{M}_{k})$ represents customer $i$’s profit contribution. By taking the derivative of $R(\mathbf{p};\mathcal{M}_{k})$ with respect to each brand’s price $p_{j}$, we characterize the optimal uniform pricing condition under a demand model $\mathcal{M}_{k}$:
Definition 1.
Profit-Maximizing Uniform Price under Model $\mathcal{M}_{k}$: Let $\alpha_{i}^{\mathcal{M}_{k}}$ denote the price sensitivity estimate for each customer $i$ under model $\mathcal{M}_{k}$. Then the optimal uniform price vector $\mathbf{p}_{*}=(p_{*,1},\ldots,p_{*,J})^{\prime}$ satisfies the following system of equations:
$p_{*,j}-m_{j}=c_{j}(\mathbf{p}_{*};\mathcal{M}_{k})+\sum_{i=1}^{N}w_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})R_{i}(\mathbf{p}_{*};\mathcal{M}_{k}),$ (22)
where $c_{j}(\mathbf{p}_{*};\mathcal{M}_{k})=-\sum_{i=1}^{N}\frac{Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})}{\sum_{i^{\prime}=1}^{N}\alpha_{i^{\prime}}^{\mathcal{M}_{k}}Pr_{i^{\prime}j}(\mathbf{p}_{*};\mathcal{M}_{k})}$ and $w_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})=\frac{\alpha_{i}^{\mathcal{M}_{k}}Pr_{ij}(\mathbf{p}_{*};\mathcal{M}_{k})}{\sum_{i^{\prime}=1}^{N}\alpha_{i^{\prime}}^{\mathcal{M}_{k}}Pr_{i^{\prime}j}(\mathbf{p}_{*};\mathcal{M}_{k})}$. Further, when we use the naive model, $\alpha_{i}^{\mathcal{M}_{N}}=\tilde{\alpha}_{0}$ and when we use our model, the $\alpha_{i}^{\mathcal{M}_{O}}=\alpha_{0}+\alpha_{2}\times\textrm{cumPercOnlineSpend}_{i}$.
This formulation gives us $J$ equations corresponding to $J$ unknown prices. We can numerically solve for the optimal price vectors based on this system of $J$ equations for a given demand model $\mathcal{M}_{k}$. See Web Appendix §$\S$§F.2 for detailed proofs and derivations.
We denote the profit-maximizing uniform prices solution for a given demand model $\mathcal{M}_{k}$ as $\mathbf{p}_{*}^{U,\mathcal{M}_{k}}$, and $k\in\{N,O\}$. Specifically, $\mathbf{p}_{*}^{U,\mathcal{M}_{N}}$ denotes the profit-maximizing uniform prices under the naive model $\mathcal{M}_{N}$ and $\mathbf{p}_{*}^{U,\mathcal{M}_{O}}$ denotes the profit-maximizing uniform prices under our model $\mathcal{M}_{O}$.
9.2.2 Personalized Pricing
We now consider a personalized pricing strategy where the firm can tailor prices to each customer based on their price elasticity through targeted coupons and email campaigns. Personalized pricing is now commonly used by firms in retail settings and has been the focus of recent research in marketing (Rossi et al. 1996; Smith et al. 2023; Jain et al. 2024).
Here, the retailer’s objective is to maximize the profit for each customer $i$:
$R_{i}(\mathbf{p}^{i};\mathcal{M}_{k})=\sum_{j=1}^{J}(p^{i}_{j}-m_{j})\times Pr_{ij}(\mathbf{p}^{i};\mathcal{M}_{k}),$ (23)
where $\mathbf{p}^{i}\in\mathcal{R}^{J\times 1}$ is the personalized price vector offered to customer $i$. As before, let $Pr_{ij}(\mathbf{p}^{i};\mathcal{M}_{k})=\frac{\exp(\psi_{ij}(\mathbf{p}^{i};\mathcal{M}_{k})}{1+\sum_{j^{\prime}=1}^{J}\exp(\psi_{ij^{\prime}}(\mathbf{p}^{i};{\mathcal{M}}_{k}))}$ denote the probability that customer $i$ chooses brand $j$ under model $\mathcal{M}_{k}$ given price $\mathbf{p}^{i}$. Following Jain et al. (2024), we show that there exists a unique globally optimal revenue and price vector characterized by a fixed-point equation and derive the optimal price vector as follows.
Definition 2.
Profit-Maximizing Personalized Price under Model $\mathcal{M}_{k}$: The optimal personalized price vector under model $\mathcal{M}_{k}$ for any customer $i$, $\mathbf{p}^{i,\mathcal{M}_{k}}_{*}:=\arg\max_{\mathbf{p}^{i}\in\mathcal{R}^{J\times 1}_{\geq 0}}R_{i}=R_{i}(\mathbf{p}^{i};\mathcal{M}_{k})$ satisfies:
$p^{i}_{*,j}-m_{j}=-\frac{1}{\alpha_{i}^{\mathcal{M}_{k}}}+R_{i},$ (24)
where $p^{i}_{*,j}$ is the optimal price for brand $j$ for customer $i$, $\alpha_{i}^{\mathcal{M}_{k}}$ is the price sensitivity estimate of customer $i$. Further, $R_{i}(\mathbf{p}^{i}_{*};\mathcal{M}_{k})$ is the solution of the fixed point equation:
$R_{i}(\mathbf{p}^{i};\mathcal{M}_{k})=-\frac{1}{\alpha_{i}^{\mathcal{M}_{k}}}\sum_{j=1}^{J}\exp(-1+\alpha_{i}^{\mathcal{M}_{k}}\cdot m_{j}+v_{ij}({\mathcal{M}_{k}}))\exp(\alpha_{i}^{{}^{\mathcal{M}_{k}}}\cdot R_{i}(\mathbf{p}^{i};\mathcal{M}_{k})),$ (25)
where $v_{ij}({\mathcal{M}_{k}})$ represents the sum of all remaining terms in the utility for customer $i$ and brand $j$ under model $\mathcal{M}_{k}$ that do not include price.
Notice that Equation (25) has a unique solution since the left-hand side is increasing and starts from zero, and the right-hand side is positive at $R_{i}(\mathbf{p}^{i};\mathcal{M}_{k})=0$ and decreasing. Following Jain et al. (2024), we use a binary search algorithm to obtain the optimal $R_{i}(\mathbf{p}^{i}_{*};\mathcal{M}_{k})$ for each individual consumer $i$ and then calculate her optimal price vector $\mathbf{p}_{*}^{i,\mathcal{M}_{k}}$. The detailed proofs for personalized pricing definitions are shown in Web Appendix §F.3. Finally, we denote the profit-maximizing personalized prices for each customer $i$ under the naive and our model as $\mathbf{p}_{*}^{i,\mathcal{M}_{N}}$ and $\mathbf{p}_{*}^{i,\mathcal{M}_{O}}$, respectively.
9.3 Profit Evaluation under Counterfactual Scenarios
We now consider four counterfactual scenarios that allow for different pricing strategies (uniform vs. personalized pricing) and whether the retailer accounts for the online adoption effect (naive vs. our model). Each scenario is characterized by its input parameters, including the individual price sensitivity parameter $\alpha_{i}$ and the mean utility function $\psi_{ij}(\mathbf{p};\mathcal{M}_{k})$ as defined in §9.2. Specifically, we consider four counterfactual scenarios, where the firm sets profit-maximizing prices:
  • $\bullet$
    Uniform Price using the Naive Model, denoted by $\mathbf{p}_{*}^{U,\mathcal{M}_{N}}$.
  • $\bullet$
    Uniform Price using Our Model, denoted by $\mathbf{p}_{*}^{U,\mathcal{M}_{O}}$.
  • $\bullet$
    Personalized Price under the Naive Model, denoted by $\mathbf{p}_{*}^{i,\mathcal{M}_{N}}$ for customer $i$.
  • $\bullet$
    Personalized Price under Our Model, denoted by $\mathbf{p}_{*}^{i,\mathcal{M}_{O}}$ for customer $i$.
The counterfactual analysis consists of two steps: (1) Optimal price calculation and (2) Profit evaluation. In the optimal price calculation step, we numerically solve for the optimal uniform and personalized price vectors under both our model and the naive model, as described in §9.2.1 and §9.2.2. Next, in the profit evaluation step, we take the optimal price from Step 1 and evaluate the counterfactual profit using Our model $\mathcal{M}_{O}$ (which we treat as the true model that accounts for the online adoption effect). This gives us the counterfactual profits for all four scenarios described above.
Next, we define how much the firm would gain if it uses our model instead of the naive model and engages in uniform pricing, as follows:
ΔGain from Uniform Pricing (%)=$\Delta\textrm{Gain from Uniform Pricing (\%)}=\frac{R(\mathbf{p}_{*}^{\mathcal{U,M}_{O}};\mathcal{M}_{O})-R(\mathbf{p}_{*}^{U,\mathcal{M}_{N}};\mathcal{M}_{O})}{R(\mathbf{p}_{*}^{U,\mathcal{M}_{N}};\mathcal{M}_{O})}$ (26)
Similarly, we can define the gains from using our model instead of the naive model for the personalized pricing case:
ΔGain from Personalized Pricing (%)=$\Delta\textrm{Gain from Personalized Pricing (\%)}=\frac{\Sigma_{i=1}^{N}R_{i}(\mathbf{p}_{*}^{i,\mathcal{M}_{O}};\mathcal{M}_{O})-\Sigma_{i=1}^{N}R_{i}(\mathbf{p}_{*}^{i,\mathcal{M}_{N}};\mathcal{M}_{O})}{\Sigma_{i=1}^{N}R_{i}(\mathbf{p}_{*}^{i,\mathcal{M}_{N}};\mathcal{M}_{O})}$ (27)
Finally, we simulate the four counterfactual scenarios using data from June 2021. We focus on the last month of the observational period to provide a clear interpretation of the firm’s monthly incremental profit and to avoid any time-specific variations across months that could complicate the interpretation. Table A21 in Web Appendix F.4 summarizes the sample sizes for each category included in this counterfactual study (i.e., customers defined in §3.2 with a purchase occasion in June 2021).
We now discuss the results of our counterfactual analysis. Under uniform pricing, we find that using our model (accounts for the online adoption effect) to optimize prices, increases monthly profit by 1.17% compared to pricing using the naive model. Further, under personalized pricing, using our model to optimize prices increases monthly profit by 4.1% compared to pricing using a naive model. Overall, these results suggest that the retailer can benefit from accounting for the effect of online adoption on consumers’ offline price sensitivity when optimizing their offline pricing decisions. Further, the magnitude of these gains is non-trivial, economically meaningful and consistent with the earlier literature on personalization (Rossi et al. 1996; Yoganarasimhan et al. 2023; Rafieian and Yoganarasimhan 2023).
Overall, our findings highlight an important yet overlooked managerial implication for multichannel retailers: retailers need to incorporate online adoption effects into offline pricing strategies, and doing so can improve retailers’ profitability in the offline channel, regardless of whether pricing is uniform or personalized. Further, the magnitude of these gains is economically meaningful. As retailers continue to expand their multichannel operations, adjusting for this effect in offline pricing becomes increasingly important.
Nevertheless, our counterfactual results should be interpreted carefully, with the appropriate caveats. First, it is possible that consumers’ likelihood of shopping offline changes under the counterfactual pricing strategy. However, consistent with the earlier literature studying brand choice within channels, (Danaher et al. 2003; Chu et al. 2008), we do not model the channel choice decision (Valentini et al. 2011), and this can potentially bias our counterfactual results. Second, we abstract away from retailer-level competition, and we do not model consumers’ choices between different retailers and how changes to offline prices may affect these choices(Bell et al. 1998). Again, this is outside of the scope of our data. Finally, our counterfactual policy estimates are based on model-based estimates rather than based on off-policy IPW evaluation (such as Smith et al. (2023)). This is because there is insufficient overlap in price distributions between monopolistic profit-maximizing prices and observed prices in the data. Nevertheless, in line with model-based evaluation similar to Rossi et al. (1996), these counterfactual analyses provide important insights into the potential gains from accounting for the effect of online adoption on offline pricing. We hope that future research can build on our findings and make further progress on these issues.
10 Conclusions
@@
The rapid growth of e-commerce has fundamentally reshaped consumer behavior, but its spillover effects on offline shopping dynamics remain understudied. In this paper, we investigate how the adoption of online shopping influences consumers’ offline price sensitivity, utilizing transaction-level data from a multichannel pet supplies retailer in Brazil. We find that consumers who adopt online shopping become more price-sensitive in their offline purchases. The relative increase in offline price elasticity, defined as the ATT relative to the pre-adoption average price elasticity of the treated group, is quite heterogeneous across categories, with a 72.7% and a 34.7% increase for dog hygiene and cat hygiene categories, and a smaller increase of 8.4% for dog food. We find no effect on cat food, presumably because it is hard to switch brands easily in this category.
From a substantive standpoint, our results contribute to the literature on price elasticity by identifying online adoption as a novel determinant of offline price sensitivity. For practitioners, our findings have significant implications on retailer offline pricing and profitability, emphasizing the need for multichannel retailers to account for the online adoption effect (i.e., heightened offline price sensitivity) in both uniform and personalized offline pricing decisions. Counterfactual simulations demonstrate that accounting for cross-channel spillovers can boost offline profits by $\approx 1-4\%$, depending on whether the retailer has the flexibility to personalize prices.
@@
Finally, our paper opens many avenues for future research. First, future research could explore whether these effects generalize to other markets (e.g., non-discretionary goods or regions with lower e-commerce penetration) and investigate mediating factors, such as the role of mobile price-checking apps or subscription models. Further, while our paper documents one aspect of consumer behavior (price sensitivity), multi-channel shopping can induce other spillover effects from one channel to another. For instance, product categories purchased offline may change after adopting online shopping (Huyghe et al. 2017), and there may be cross-media effects between channels (Joo et al. 2014; Du et al. 2019). Future research can investigate other such spillover effects and their effect on cross-channel purchases. Ultimately, as retail continues to evolve toward omnichannel integration, understanding these interdependencies will be critical for both academics and firms aiming to thrive in a digitally transformed landscape.
@@ A.1.1 Selection of Brands and Stock Keeping Units (SKUs)
For each product category (i.e., dry dog food, dry cat food, dog hygiene and cat hygiene), we focus on a single pack size that accounts for the largest sales revenue and offers sufficient brand representation. Our goal is to study consumer brand choice and price elasticity, so consistent with previous brand choice literature (Allenby and Rossi 1998; Kim et al. 2002; Chintagunta et al. 2005; Dubé et al. 2010) we focus on a market (i.e., single most popular pack size within each category) that initially (as of January 1st, 2019 - June 30th, 2019) features at least three brands. We exclude any market with fewer than three brands or with one brand contributing more than 60% of total sales. Applying these criteria yields the top one selling pack sizes for each category during the observation period: 15KG adult daily dry dog food, 10.1KG adult daily dry cat food, and an 80cm $\times$ 60cm dog hygiene mat containing 30 pieces.
@@
After defining the market with the top one selling pack size for each category, we rank brands by offline sales in descending order and include those that collectively account for at least 80% of the sales, following earlier literature (Gordon et al. 2013). The remaining brands are grouped into a single ‘Other Brand’ for each product category, ensuring broad market coverage.

### v2

@@ 8.2.3 Alternative Control Groups
[…]
8.2.4 Alternative Definition of Purchase Occasions
In §3.2, we define a purchase occasion as any customer-month during which we observe any offline interaction with the retailer (similar to a week-store visit in Gordon et al., (2013)). In this section, we consider an alternative definition. Specifically, if a customer visits the retailer (i.e., either in-store or online) during a month and purchases in any category (not limited to the categories we focus on in this study), that customer-month is included as a purchase occasion for all four categories. If a customer makes no purchases in any category during the month, that customer-month is not included as a purchase occasion. In other words, a purchase occasion represents any customer-month during which we observe any interaction with the retailer, either offline or online.
The detailed demand model estimates and ATT estimation results are presented in Web Appendix §E.6. Overall, our conclusions remain similar to those from our main specification, i.e., that consumers become more offline price-sensitive after adopting online shopping.
9 Managerial Implications and Conclusions
@@
The rapid growth of e-commerce has fundamentally reshaped consumer behavior, but its spillover effects on offline shopping dynamics remain understudied. In this paper, we investigate how the adoption of online shopping using a retailer’s e-commerce channels influences consumers’ offline price sensitivity when shopping at that retailer’s physical stores, utilizing transaction-level data from a multichannel pet supplies retailer in Brazil. We find that consumers who adopt online shopping become more price-sensitive in their offline purchases. The relative increase in offline price elasticity, defined as the ATT relative to the pre-adoption average price elasticity of the treated group, is quite heterogeneous across categories, with a 72.7% and a 34.7% increase for dog hygiene and cat hygiene categories, and a smaller increase of 8.4% for dog food. We find no effect on cat food, presumably because it is hard to switch brands easily in this category.
From a substantive standpoint, our results contribute to the literature on price elasticity by identifying online adoption as a novel determinant of offline price sensitivity. This factor has so far been ignored in the demand estimation literature, which usually only accounts for factors like consumer demographics and macro-economic shocks to affect consumer price sensitivity. From a managerial standpoint, our findings have implications for retailer offline pricing and profitability, emphasizing the need for multichannel retailers to account for the online adoption effect (i.e., heightened offline price sensitivity when consumers adopt their online channels) in their pricing decisions. Further, our approach can be used to personalize prices depending on customers’ channel adoption status. While existing literature has largely focused on pricing and personalization based on customer demographics and prior purchase behavior (Rossi and Allenby, 1993; Chen et al., 2001; Jiang et al., 2021; Jain et al., 2024), our work suggests that incorporating cross-channel effects into measures of price elasticity and pricing strategy can further enhance profits.
@@
Finally, our paper opens many avenues for future research. First, future research could explore whether these effects generalize to other markets (e.g., non-discretionary goods or regions with lower e-commerce penetration) and investigate mediating factors, such as the role of mobile price-checking apps or subscription models. Further, while our paper documents one aspect of consumer behavior (price sensitivity), multi-channel shopping can induce other spillover effects from one channel to another. For instance, product categories purchased offline may change after adopting online shopping (Huyghe et al., 2017), and there may be cross-media effects between channels (Joo et al., 2014; Du et al., 2019). Future research can investigate other such spillover effects and their effect on cross-channel purchases. Ultimately, as retail continues to evolve toward omnichannel integration, understanding these interdependencies will be critical for both academics and firms aiming to thrive in a digitally transformed landscape.
@@ A.1.1 Selection of Brands and Stock Keeping Units (SKUs)
For each product category (i.e., dry dog food, dry cat food, dog hygiene and cat hygiene), we focus on a single pack size that accounts for the largest sales revenue and offers sufficient brand representation. Our goal is to study consumer brand choice and price elasticity, so consistent with previous brand choice literature (Allenby and Rossi, 1998; Kim et al., 2002; Chintagunta et al., 2005; Dubé et al., 2010) we focus on a market (i.e., single most popular pack size within each category) that initially (as of January 1st, 2019 - June 30th, 2019) features at least three brands. We exclude any market with fewer than three brands or with one brand contributing more than 60% of total sales. Applying these criteria yields the top one selling pack sizes for each category during the observation period: 15KG adult daily dry dog food, 10.1KG adult daily dry cat food, and an 80cm $\times$ 60cm dog hygiene mat containing 30 pieces.
@@
After defining the market with the top one selling pack size for each category, we rank brands by offline sales in descending order and include those that collectively account for at least 80% of the sales, following earlier literature (Gordon et al., 2013). The remaining brands are grouped into a single ‘Other Brand’ for each product category, ensuring broad market coverage.